A bookstore sells paperback books for $5.95 each and hardcover books for $19.75 each. Last Tuesday, 86 books were sold for a total of $939.50. How many of each paperback and hardcover books were sold?
Answer by macston(5194)

.
P=number of paperbacks; H=number of hardcover books
.
P+H=86
P=86-H . Use this to substitute for P below.
.
$5.95P+$19.75H=$939.50 . Substitute for P from above.
$5.95(86-H)=$19.75H=$939.50
$511.70-$5.95H+$19.75H=$939.50
$13.80H=$427.80
H=31
.
ANSWER 1: 31 hardcover books were sold.
.
P=86-H=86-31=55
.
ANSWER 2: 55 paperback books were sold.
.
CHECK:
$5.95P+$19.75H=$939.50
$5.95(55)+$19.75(31)=$939.50
$327.25+$612.25=$939.50
$939.50=$939.50
.